OpenStudy (anonymous):

A town's population has been growing linearly. In 2003, the population was 45000, and the population has been growing by 2200 people each year. Write an equation for the population x years after 2003. P=

OpenStudy (anonymous):

@lucaz

OpenStudy (anonymous):

@EducateMe

OpenStudy (kirbykirby):

y=mx + b Well you can figure out the intercept "b".. this is just 45000. Now, the slope "m" is just the change of growth every year.

OpenStudy (anonymous):

I just don't know how many years you want this equation to represent?

OpenStudy (kirbykirby):

Well, you could say like \(y = 2200x + 45000, \text{for } x\in \{1, 2, ..., 197\}\)

OpenStudy (anonymous):

true

OpenStudy (anonymous):

But it would be easier to solve with a full equation, but that's fine.

OpenStudy (kirbykirby):

here, x=1 would represent 2004 (since they ask for after 2003), so 2200 - 2004 = 196, but you add one to count the year 2004 If you're not sure why we add 1, think of the number of years between 2004 and 2006. That's 2004, 2005 and 2006, 3 years. But 2006 - 2004 = 2.

OpenStudy (kirbykirby):

What do you mean by a full equation?